OpenStudy (anonymous):

Which of the following is the equation of a line that passes through the points (0,6) and (2,10)?

OpenStudy (kidthatbro8):

i'm not aware of your answer choices, but here, let me offer some help; http://web.geogebra.org/app/# that site there ^ type in your equations and place your points.

OpenStudy (anonymous):

Oh my bad i forgot.. here a. y = -2x - 6 b. y = 2x - 6 c. y = 2x + 6 d. y = -2x + 6

OpenStudy (aripotta):

find the slope

OpenStudy (aripotta):

\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (kidthatbro8):

the website can also help, but i must agree too with @aripotta

OpenStudy (aripotta):

they gave you the y-intercept as one of the points. all you have to do is find the slope

OpenStudy (aripotta):

then you can plug in the slop and y-intercept into the formula y = mx + b, where m is the slope and b is the y-intercept